Based on the BOM shown below, how many of part J will be needed if 40 units of end item A are needed? Managers have decided to outsource part G. Revised the BOM for end item A, assuming that item G is now purchase from a supplier. How many of part J will now be needed if 40 units of end item A are needed? Show all work.

em A, assuming that item G is now purchased from a suppliet. Hw maty of ant t now be nceded if 40 usits of end item A are needed)7 6同 1 wee 2 weeks 1 ek 1闪 t woekTwek s product stracture tree for the baking pan using the BOM shown below, If as to make 100 baking pams, how many handles are needed? How many t sets are nceded? Baking Pan
Answer #1

1) In the product tree, we see that A takes 2 Bs, 1C, 3Ds and 1E

Each B takes 2Fs and 2Gs, Each D takes 5Gs and Each E takes 4Js

Each G takes 2Is and 2Js

Bs needed = 2*A = 2*40 = 80

Ds needed = 1*A = 1*40 = 40

Es needed = 1*A = 1*40 = 40

Gs needed = 2*B + 5*D = 2*80 + 5*40 = 360

Js needed = 2*G + 4*E = 2*360 + 4*40 = 880

2) If G is purchased from a supplier,

Js needed = 4*E = 4*40 = 160